A sample of argon has a volume of 1.2 L at STP. If the temperature is increased to 22°C and the pressure is lowered to 0.80 atm, what will the new volume be, in L?
1
Expert's answer
2020-11-18T13:44:48-0500

At STP, Temperature =0°C(273K) and pressure =1atm

Using the general gas equation;

V1×P1/T1=V2×P2/T2

V1=1.2L P1=1atm T1=273k

V2= ? P2=0.8atm T2=295k


Making V2 subject of the formula

V2=V1×P1×T2/P2×T1

V2= 1.2×1×295/0.8×273

V2=1.62L